OpenStudy (howard-wolowitz):

Find the sum of the first 25 terms of a geometric sequence where the first term is 177,147 and the common ratio is -1/3

OpenStudy (howard-wolowitz):

A. 132,860.1667 B. 132,860.25

OpenStudy (howard-wolowitz):

C. 132,860

OpenStudy (howard-wolowitz):

D. 132,860.5

OpenStudy (howard-wolowitz):

@freckles

OpenStudy (howard-wolowitz):

@Michele_Laino

OpenStudy (howard-wolowitz):

HEy freckls!

OpenStudy (howard-wolowitz):

MY TWO FRIENDS!!

OpenStudy (freckles):

\[a_1=177,147 \\ r=\frac{-1}{3} \text{ and you want to add all the way up the n=25th term }\\ \sum_{i=1}^{n}a_i=a \cdot \frac{1-r^n}{1-r}\] all this info is given just plug in

OpenStudy (freckles):

a is a_1 by the way

OpenStudy (michele_laino):

please try this formula: \[{S_n} = {a_1}\frac{{1 - {q^n}}}{{1 - q}}\]

OpenStudy (howard-wolowitz):

Alright so I did that and got either C or B

OpenStudy (michele_laino):

where n= 25 and q=r =-1/3

OpenStudy (howard-wolowitz):

ok so use that to do what subsuite

OpenStudy (freckles):

you can use a calculator right?

OpenStudy (howard-wolowitz):

no

OpenStudy (howard-wolowitz):

can you though

OpenStudy (michele_laino):

hint: here is your substitution: \[{S_n} = 177147\frac{{1 - {{\left( { - 1/3} \right)}^{25}}}}{{1 - \left( { - 1/3} \right)}}\]

OpenStudy (freckles):

what? so you are really expected to do 3^(25) in your head?

OpenStudy (freckles):

Or on paper.

OpenStudy (howard-wolowitz):

alright I'll use one

OpenStudy (freckles):

ok so you are going to input: 177147 times (1-(-1/3)^(25))/(1-(-1/3))

OpenStudy (howard-wolowitz):

i got 847288609443

OpenStudy (howard-wolowitz):

i did what you said

OpenStudy (freckles):

try it again I don't think 177147 times (1-(-1/3)^(25))/(1-(-1/3)) equals that

OpenStudy (howard-wolowitz):

3^25 is though

OpenStudy (freckles):

ok i didn't ask you to do that i was just kinda like wow your teach expects you to do that without calculator

OpenStudy (howard-wolowitz):

oh ok sorry my bad

OpenStudy (michele_laino):

please use Windows calculator @Howard-Wolowitz

OpenStudy (howard-wolowitz):

ok im doing it one sec

OpenStudy (howard-wolowitz):

dang i got a big number

OpenStudy (howard-wolowitz):

−125099989649180177147

OpenStudy (howard-wolowitz):

-125099989649180/177147

OpenStudy (freckles):

177,147*(1-(-1/3)^(25))/(1-(-1/3)) hmmm I don;t you input this because somehow you put 177147 on bottom

OpenStudy (howard-wolowitz):

so off of that site im going with C

OpenStudy (freckles):

I think another choice is closer

OpenStudy (michele_laino):

I got a different option

OpenStudy (howard-wolowitz):

do i use one of the .25 options from the site

OpenStudy (freckles):

that choice seems to be a lot closer

OpenStudy (howard-wolowitz):

closer?

OpenStudy (freckles):

nearer to the value on that page I sent you

OpenStudy (freckles):

it said something like 132860.2500001568

OpenStudy (freckles):

132860 is close but 132860.25 is a lot closer

OpenStudy (michele_laino):

using Windows calc I got this value: 132860,24999984319363140342327119

OpenStudy (freckles):

and @Michele_Laino value he got is still closer to 132860.25 than it is to 132860

OpenStudy (howard-wolowitz):

ok so i got what he got then?

OpenStudy (freckles):

which is what?

OpenStudy (howard-wolowitz):

.25

OpenStudy (michele_laino):

that's right!

OpenStudy (freckles):

well you mean the 132860.25 and yes that is what i got above

OpenStudy (howard-wolowitz):

well i thought u meant i was close not right

OpenStudy (freckles):

no I was saying out of all the choices 132860.25 is the closest

OpenStudy (michele_laino):

please I think that you have to round off your result @Howard-Wolowitz

OpenStudy (howard-wolowitz):

i gotcha

OpenStudy (howard-wolowitz):

ok i got it

OpenStudy (howard-wolowitz):

.5

OpenStudy (howard-wolowitz):

is .5 right

OpenStudy (howard-wolowitz):

@freckles